I must solve the equation using the substitution method.
2x+y=3
y=5x-11
I worked the problem, but i dont know if it is correct...i solved as follows..
2x+5x-11=3
7x-11=3
+11 +11
7x = 14
-- --
7 7
x = 2
Then I solve for y as follows:
y=5(2)-11
y=10-11
y = -1
So, my answer is (2,-1)
Is this correct??

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

I must solve the equation using the substitution method.
2x+y=3
y=5x-11
-----------
Substitute for "y" and solve for "x":
2x+ 5x-11 = 3
7x = 14
x = 2
----
Substitute into y = 5x-11 to solve for "x":
y = 5*2-11 = -1
-----------------------
Solution: (2,-1)
========================
Cheers,
Stan H.
==========

Answer by Alan3354(69443)

2x+y=3
y=5x-11
-----------
So, my answer is (2,-1)
Is this correct??
---------------
As a check, see if your answer fits both equations.
2x+y = 3
2*2 + (-1) =? 3
3 = 3 That one's good.
-------------
y=5x-11
-1 =? 5*2 - 11
-1 = -1 both fit, so you're correct.